# Simultaneous Khintchine theorem on manifolds in positive characteristics: convergence case

### v1

@@ Abstract.
In this article, we prove the convergence case of Khintchine’s theorem for analytic nonplanar manifolds over local fields of positive characteristic, in the setting of simultaneous Diophantine approximation. Our approach is based on the method of counting rational points near manifolds developed by Beresnevich and Yang [BY23]. The results obtained here extend the work of Beresnevich and Yang [BY23], and Beresnevich and Datta [BD25], to the function field setting. In the course of the proof, we also establish several new results in the geometry of numbers over function fields, which are of independent interest.
@@ 1. Introduction
Diophantine approximation is a central topic in number theory, which deals with the effective density of rational numbers within real numbers and its higher-dimensional analogue. The Khintchine-Groshev theorem is a foundational result in metric Diophantine approximation. Let $\psi:(0,\infty)\rightarrow(0,1)$ be a approximating function, i.e., a non-increasing function such that $\psi(x)\rightarrow 0$ as $x\rightarrow\infty.$ For a fixed $\bm{\theta}\in\mathbb{R}^{n}$ with $n\in\mathbb{N}$, the following set is an object of interest for a long time
$\mathcal{S}_{n}^{\bm{\theta}}(\psi):=\left\{\mathbf{y}\in\mathbb{R}^{n}:\left\|\mathbf{y}-\frac{\mathbf{p}+\bm{\theta}}{q}\right\|<\frac{\psi(q)}{q}\,\,\text{for i.m. }(\mathbf{p},q)\in\mathbb{Z}^{n}\times\mathbb{N}\right\},$
where $\|\cdot\|$ denotes the sup norm and ‘i.m’ reads as ‘infinitely many’. The points $\mathbf{y}$ lying in $\mathcal{S}_{n}^{\bm{\theta}}(\psi)$ are usually referred to as $(\psi,\bm{\theta})$-approximable. The case $\bm{\theta}=\mathbf{0}$ is called homogeneous setting, and the points of $\mathcal{S}_{n}^{\mathbf{0}}(\psi)$ are called $\psi$-approximable. The form of approximation described above concerns how well one can approximate points of $\mathbb{R}^{n}$ by rational points, and it is commonly referred to as the simultaneous form of approximation. In contrast, there is another widely studied type of approximation, where one asks how close a point in $\mathbb{R}^{n}$ is to a rational hyperplane; this is known as the dual form of approximation. To begin with, we recall the inhomogeneous version of classical Khintchine’s theorem (see [Khi26, Gro38] and [AR23, §1.1]).
@@
$\mathcal{S}_{n}^{\bm{\theta}}(\psi)=\begin{cases}\text{Lebesgue null}&\text{if }\displaystyle\sum_{q=1}^{\infty}\psi(q)^{n}<\infty\\ \text{Lebesgue full}&\text{if }\displaystyle\sum_{q=1}^{\infty}\psi(q)^{n}=\infty.\end{cases}$ (1.1)
The situation becomes more delicate when examining the extent to which embedded submanifolds of $\mathbb{R}^{n}$ inherit the Diophantine properties prevalent in $\mathbb{R}^{n}$. Establishing an analogue of Khintchine’s theorem for manifolds is closely linked to the challenging problem of counting rational points near the manifold (for example, see [Ber02, Ber12, BVVZ21, SST23, BD25]). For analytic non-degenerate manifolds of dimension $d\geq 2$, the divergence case was established by Beresnevich [Ber12], and later extended to non-analytic curves in [BVVZ21], both employing the ubiquity framework alongside Kleinbock–Margulis’s quantitative non-divergence estimates [KM98]. For non-degenerate manifolds in $\mathbb{R}^{n}$, the convergence case was resolved by Beresnevich and Yang [BY23] (see also [SST23]), and Khintchine’s theorem in full generality for simultaneous approximation was recently settled by Beresnevich and Datta [BD25]. More recently, Datta [Dat24] extended the result in Beresnevich and Yang by proving a quantitative simultaneous Khintchine theorem over manifolds in $\mathbb{R}^{n}$.
Unlike the simultaneous form, the dual form of Diophantine approximation has long been well understood, beginning with Kleinbock and Margulis’s resolution of the Baker-Sprindžuk conjecture [KM98]. A complete dual Khintchine theorem for non-degenerate manifolds was later established in both homogeneous [BKM01, BBKM02] and inhomogeneous settings [BBV13].
@@
In the metric theory over manifolds in this setting, the analogues of the Baker-Sprindžuk conjecture had been proved by Ghosh [Gho07]. Subsequently, Ganguly and Ghosh proved the inhomogeneous Sprindžuk conjecture in [GG20]. Recently, Das and Ganguly proved a complete inhomogeneous Khintchine-Groshev type theorem for nonplanar manifolds over function fields in the dual setting [DG22]. For additional results and recent developments on Diophantine approximation in function fields, see [GG19, Gan22, DG24, Ban25, AK25, DX24, KLP23].
In this paper, we prove the convergence case of Khintchine’s theorem for analytic non-planar manifolds over local fields of positive characteristic, extending the earlier works of Beresnevich and Yang [BY23], and Beresnevich and Datta [BD25] to the setting of function fields. Although our approach is based on the technique of Beresnevich and Yang [BY23] for counting rational points near manifolds, the positive characteristic setup poses many challenges, which we have explained in §1.4. Apart from this, we have also proved some new results in the geometry of numbers over function fields (see §3 and §6 ), which are of independent interest. We begin by introducing the function field framework.
@@ 1.1. The Function Field Setting
The completion of $\mathbb{F}_{q}(T)$ with respect to this absolute value is the field $\mathbb{F}_{q}((T^{-1}))$, i.e., the field of Laurent series in $T^{-1}$ over the finite field $\mathbb{F}_{q}.$ The absolute value on $\mathbb{F}_{q}((T^{-1}))$, which we again denote by $|\cdot|$, is defined as follows. For $a\in\mathbb{F}_{q}((T^{-1}))$, if $a=0$, then $|a|=0$. Otherwise, $a$ can be expressed uniquely as a Laurent series,
@@
where $k_{0}\in\mathbb{Z}$, each $a_{k}\in\mathbb{F}_{q}$, and the leading coefficient $a_{k_{0}}\neq 0$. We define the degree of $a$ by $\deg(a):=k_{0}$, and set $|a|:=q^{\deg(a)}$. This clearly extends the absolute value $|\cdot|$ on $\mathbb{F}_{q}(T)$ to $\mathbb{F}_{q}((T^{-1}))$, and the extension remains non-archimedean and discrete. Consequently, $\mathbb{F}_{q}((T^{-1}))$ is a complete and separable metric space, which is ultrametric and, hence, totally disconnected. It is worth noting that any local field of positive characteristic is isomorphic to some $\mathbb{F}_{q}((T^{-1}))$.
@@
Throughout the paper, we let $\mathcal{R}$ and $\mathcal{K}$ denote, respectively, the polynomial ring $\mathbb{F}_{q}[T]$ and the field of Laurent series $\mathbb{F}_{q}((T^{-1})).$ We denote by $\mathcal{O}_{\mathcal{K}}$, the ring of integers of $\mathcal{K}$ , defined by $\mathcal{O}_{\mathcal{K}}:=\{x\in\mathcal{K}:|x|\leq 1\}$. Let $\mathcal{L}_{1}$ denote the Haar measure on $\mathcal{K}$, normalized so that $\mathcal{L}_{1}\left(\mathcal{O}_{\mathcal{K}}\right)=1$.
@@
Note that since the root of the equation $x^{\ell}=T$ is not in $\mathcal{K}$ for every $\ell\geq 2$, the degree of the algebraic closure of $\mathcal{K}$ over $\mathcal{K}$ is $\infty$. Nevertheless, one can discuss finite extensions, such as the extension of the function field $\mathcal{K}$ obtained by adjoining the $\ell$-th root of $T$, and we denote this extension by $\mathcal{K}_{\ell}$. Let the polynomial ring in $\mathcal{K}_{\ell}$ be denoted by $\widetilde{\mathcal{R}}$, i.e., $\widetilde{\mathcal{R}}:=\mathbb{F}_{q}[T^{1/\ell}]\subset\mathcal{K}_{\ell}$. So we have the natural inclusion $\mathcal{R}\subset\widetilde{\mathcal{R}}\subset\mathcal{K}_{\ell}.$ It is worth noting that $\widetilde{\mathcal{R}}$ is a lattice in $\mathcal{K}_{\ell}$, while $\mathcal{R}$ is only a discrete subgroup when viewed as a subset of $\mathcal{K}_{\ell}$. Given $n\geq 2$, we equip $\mathcal{K}^{n}$ with the sup norm defined by
$\|\mathbf{x}\|=\max_{1\leq i\leq n}|x_{i}|,\quad\text{for}\,\,\mathbf{x}=(x_{1},\dots,x_{n})\in\mathcal{K}^{n}.$
We also equip $\mathcal{K}_{\ell}^{n}$ with an analogous sup norm. Throughout, $\mathcal{L}_{n}$ denotes the product Haar measure on $\mathcal{K}^{n}$. From now on, we always assume $n:=d+m,$ where $d,m\in\mathbb{N}$.
@@
Given a non-increasing function $\psi:(0,\infty)\to(0,1)$ and a vector $\bm{\Theta}\in\mathcal{K}^{n}$, we define the set of all $(\psi,\bm{\Theta})$-approximable vectors in $\mathcal{K}^{n}$ by
@@
$\mathcal{S}_{n}^{\bm{\Theta}}(\psi):=\left\{\mathbf{x}\in\mathcal{K}^{n}:\left\|\mathbf{x}-\frac{\mathbf{P}+\bm{\Theta}}{Q}\right\|<\frac{\psi(|Q|)}{|Q|}\quad\text{for infinitely many }(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\})\right\}.$
@@
In the special case when $\bm{\Theta}=\mathbf{0}$, the set reduces to the set of $\psi$-approximable vectors, which we denote by $\mathcal{S}_{n}({\psi}):=\mathcal{S}_{n}^{\mathbf{0}}(\psi).$
@@
Without loss of generality, we consider the manifold $\mathscr{M}$ defined by the map $\mathbf{f}:U\subset\mathcal{K}^{d}\to\mathcal{K}^{n}$, where
$\mathbf{f}:=(x_{1},\ldots,x_{d},f_{1}(\mathbf{x}),\ldots,f_{m}(\mathbf{x}))=(\mathbf{x},\bm{f}(\mathbf{x})),$
@@
with $d=\dim(\mathscr{M})$, $m=\operatorname{\text{codim}}(\mathscr{M})$, and $U\subseteq\mathcal{K}^{d}$ is an open set. We also assume that the defining map $\mathbf{f}$ is analytic (see section˜2 for the definition).
@@ 1.2. Main Results
Throughout the paper, we make the following assumptions:
@@
    The map $\bm{f}=(f_{1},\dots,f_{m}):U\subseteq\mathcal{K}^{d}\rightarrow\mathcal{K}^{m}$ is an analytic map that can be analytically extended to the boundary of $U$.
@@
  3. (III)
    We always assume $\psi:(0,\infty)\rightarrow(0,1)$ to be an approximating function, i.e., $\psi$ is non-increasing and $\psi(x)\rightarrow 0$ as $x\rightarrow\infty.$ Furthermore, without loss of generality, we can always assume that the range of $\psi$ is contained in $\{q^{-k}:k\in\mathbb{N}\}.$
Without loss of generality, for the sake of simplification, we further make some boundedness assumptions:
  1. (IV)
    $\|\bm{f}(\mathbf{x})\|\leq 1,$ $\|\nabla\bm{f}(\mathbf{x})\|\leq 1,$ and $|\bar{\Phi}_{\beta}\bm{f}(\mathbf{y_{1}},\mathbf{y_{2}},\mathbf{y_{3}})|\leq M$ for any second-order difference quotient $\Phi_{\beta}$ and $\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}},\mathbf{y_{3}}\in U$ (see section˜2 for the definitions).
@@
Let $n\geq 2$, let $\bm{\Theta}\in\mathcal{K}^{n}$, and let $\mathscr{M}\subset\mathcal{K}^{n}$ be an analytic submanifold satisfying (I), (II), and (IV). Suppose that the approximating function $\psi:(0,\infty)\to(0,1)$ satisfies
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\psi(|Q|)^{n}<\infty.$
Then almost every point of $\mathscr{M}$ is not $(\psi,\bm{\Theta})$-approximable, i.e.,
$\mathcal{L}_{d}\left(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi))\right)=0.$
We now recall the definition of Hausdorff $\sigma$-measure. For $\sigma\geq 0$, the Hausdorff $\sigma$-measure of a set $A\subset\mathcal{K}^{n}$ is defined as
$\mathcal{H}^{\sigma}(A):=\lim_{\rho\to 0^{+}}\inf\left\{\sum_{i=1}^{\infty}\operatorname{diam}(A_{i})^{\sigma}:A\subset\bigcup_{i=1}^{\infty}A_{i}\quad\text{ and }\quad\operatorname{diam}(A_{i})<\rho,\,\,\forall\,\,i\right\},$
where $\operatorname{diam}(A)=\sup\{\|\mathbf{u}-\mathbf{v}\|:\mathbf{u},\mathbf{v}\in A\}$ denotes the diameter of a set. Now we state the Hausdorff measure analogue of Theorem 1.2.

### v2

# arXiv:2511.01991v2[math.NT] 02 Mar 2026

@@ Abstract.
We prove the convergence case of Khintchine’s theorem, with general approximation functions $\psi$ that are not necessarily monotonic, for analytic nonplanar manifolds over local fields of positive characteristic. Our approach is based on the method of counting rational points near manifolds developed by Beresnevich and Yang [BY23]. To address the scenario where the given $\psi$ is not monotonic, we extend our function field by adjoining an appropriate root. Additionally, in the course of the proof, we establish several new results in the geometry of numbers over function fields, which are of independent interest.
@@ 1. Introduction
Diophantine approximation is a central topic in number theory, which deals with the effective density of rational numbers within real numbers and its higher-dimensional analogue. The Khintchine-Groshev theorem is a foundational result in metric Diophantine approximation. Let $\psi:(0,\infty)\longrightarrow(0,1)$ be a approximating function, i.e., $\psi(x)\rightarrow 0$ as $x\rightarrow\infty$. For a fixed $\bm{\theta}\in\mathbb{R}^{n}$ with $n\in\mathbb{N}$, the following set has been an object of interest for a long time
$\mathcal{S}_{n}^{\bm{\theta}}(\psi){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\left\{\mathbf{y}\in\mathbb{R}^{n}:\left\|\mathbf{y}-\frac{\mathbf{p}+\bm{\theta}}{q}\right\|<\frac{\psi(q)}{q}\,\,\text{for i.m. }(\mathbf{p},q)\in\mathbb{Z}^{n}\times\mathbb{N}\right\},$
where $\|\cdot\|$ denotes the sup norm and ‘i.m’ reads as ‘infinitely many’. The points $\mathbf{y}$ lying in $\mathcal{S}_{n}^{\bm{\theta}}(\psi)$ are usually referred to as $(\psi,\bm{\theta})$-approximable. The case $\bm{\theta}=\mathbf{0}$ is called the homogeneous setting, and the points of $\mathcal{S}_{n}^{\mathbf{0}}(\psi)$ are referred to as $\psi$-approximable. The form of approximation described above concerns how well one can approximate points of $\mathbb{R}^{n}$ by rational points, and it is commonly referred to as the simultaneous form of approximation. In contrast, another widely studied type of approximation asks how close a point in $\mathbb{R}^{n}$ is to a rational hyperplane; this is known as the dual form of approximation. To begin with, we recall the inhomogeneous version of classical Khintchine’s theorem (see [Khi26, Gro38] and [AR23, §1.1]).
@@
[…]
It is important to note that, while one does not need any sort of monotonicity assumption to prove the convergence part, the question of whether the emphasized monotonicity assumption can be dropped in the divergence case has attracted many mathematicians over the last six or seven decades at least. For $n=1$, this assumption is crucial, owing to the counterexample constructed by Duffin and Schaeffer [DS41]. In fact, it has been shown in [Ram17] that, when $n=1$, for no $\theta\in\mathbb{R}$ one can remove the monotonicity. In a recent paper [Yu21], Han Yu showed that the monotonicity assumption can be removed for all $n\geq 3$. In the case of $n=2$, Allen and Ramírez were able to remove the same at the cost of an extra divergence of a measure sum [AR23, Theorem 3].
The situation becomes more delicate when examining the extent to which embedded submanifolds of $\mathbb{R}^{n}$ inherit the Diophantine properties prevalent in $\mathbb{R}^{n}$. Establishing an analogue of Khintchine’s theorem for manifolds is closely linked to the challenging problem of counting rational points near the manifold (for example, see [Ber02, Ber12, BVVZ21, SST23, BD25]). For analytic non-degenerate manifolds of dimension $d\geq 2$, the divergence case was established by Beresnevich [Ber12], and later extended to non-analytic curves in [BVVZ21], both employing the ubiquity framework alongside Kleinbock–Margulis’s quantitative non-divergence estimates [KM98]. For non-degenerate manifolds in $\mathbb{R}^{n}$, the convergence case was resolved by Beresnevich and Yang [BY23] (see also [SST23]), and Khintchine’s theorem in full generality for simultaneous approximation was recently settled by Beresnevich and Datta [BD25]. More recently, Datta [Dat24] extended the result in Beresnevich and Yang by proving a quantitative simultaneous Khintchine theorem over manifolds in $\mathbb{R}^{n}$. We ask the reader to note that all have assumed the approximation function is non-increasing, even for the convergence case.
On a related note, we remind the reader that, unlike the simultaneous form, the dual form of Diophantine approximation is well understood. Beginning with Kleinbock and Margulis’s resolution of the Baker-Sprindžuk conjecture [KM98]. A complete dual Khintchine theorem for non-degenerate manifolds was later established in both homogeneous [BKM01, BBKM02] and inhomogeneous settings [BBV13].
@@
In the metric theory over manifolds in this setting, the analogues of the Baker-Sprindžuk conjecture had been proved by Ghosh [Gho07]. Subsequently, Ganguly and Ghosh proved the inhomogeneous Sprindžuk conjecture in [GG20]. Recently, Das and Ganguly proved a complete inhomogeneous Khintchine-Groshev type theorem for nonplanar manifolds over function fields in the dual setting [DG25]. For additional results and recent developments on Diophantine approximation in function fields, see [GG19, Gan22, DG24, Ban25, AK25, DX24, KLP23].
In this paper, we prove the convergence case of Khintchine’s theorem for analytic non-planar manifolds over local fields of positive characteristic without the monotonicity assumption. While this certainly extends the earlier works of Beresnevich and Yang [BY23], and Beresnevich and Datta [BD25] to the setting of function fields, we believe that the fact that our result applies to general approximation functions that need not be monotonic makes it distinguishable from its real counterpart. Although our approach is based on the technique of Beresnevich and Yang [BY23] for counting rational points near manifolds, the positive characteristic setup and the absence of the monotonicity necessitate proper adaptation, which we explain in §1.4. On our way, we also prove some new results in the geometry of numbers over function fields (see §3 and §6), which are of independent interest. We begin by introducing the function field framework.
@@ 1.1. The Function Field Setting
The completion of $\mathbb{F}_{q}(T)$ with respect to this absolute value is $\mathbb{F}_{q}((T^{-1}))$, i.e., the field of Laurent series in $T^{-1}$ over the finite field $\mathbb{F}_{q}.$ The absolute value on $\mathbb{F}_{q}((T^{-1}))$, which we again denote by $|\cdot|$, is defined as follows. For $a\in\mathbb{F}_{q}((T^{-1}))$, if $a=0$ then $|a|=0$. Otherwise, $a$ can be expressed uniquely as a Laurent series,
@@
where $k_{0}\in\mathbb{Z}$, each $a_{k}\in\mathbb{F}_{q}$, and the leading coefficient $a_{k_{0}}\neq 0$. We define the degree of $a$ by $\deg(a){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}k_{0}$, and set $|a|{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}q^{\deg(a)}$. This clearly extends the absolute value $|\cdot|$ on $\mathbb{F}_{q}(T)$ to $\mathbb{F}_{q}((T^{-1}))$, and the extension remains non-archimedean and discrete. Consequently, $\mathbb{F}_{q}((T^{-1}))$ is a complete and separable metric space that is ultrametric and, hence, totally disconnected. It is worth noting that any local field of positive characteristic is isomorphic to some $\mathbb{F}_{q}((T^{-1}))$ ([Wei95, Part I, Chapter I, Theorem 8]).
@@
Throughout the paper, we let $\mathcal{R}$ and $\mathcal{K}$ denote, respectively, the polynomial ring $\mathbb{F}_{q}[T]$ and the field of Laurent series $\mathbb{F}_{q}((T^{-1})).$ We denote by $\mathcal{O}_{\mathcal{K}}$ the ring of integers of $\mathcal{K}$, defined by $\mathcal{O}_{\mathcal{K}}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\{x\in\mathcal{K}:|x|\leq 1\}$. Let $\mathcal{L}_{1}$ denote the Haar measure on $\mathcal{K}$, normalized so that $\mathcal{L}_{1}\left(\mathcal{O}_{\mathcal{K}}\right)=1$. For any $n\in\mathbb{N}$, we equip $\mathcal{K}^{n}$ with the sup norm, which is defined as follows:
$\|\mathbf{x}\|=\max_{1\leq i\leq n}|x_{i}|,\quad\text{for}\,\,\mathbf{x}=(x_{1},\dots,x_{n})\in\mathcal{K}^{n},$
and we equip it with the product Haar measure, denoted by $\mathcal{L}_{n}$.
@@
The field $\mathcal{K}=\mathbb{F}_{q}((T^{-1}))$ is not algebraically closed. For every integer $\ell\geq 2$, the equation $x^{\ell}=T$ has no solution in $\mathcal{K}$. Consequently, the algebraic closure of $\mathcal{K}$ is an infinite algebraic extension of $\mathcal{K}$. (see [Ked01] for more information about the algebraic closure of $\mathcal{K}$). Nevertheless, one can discuss finite extensions, such as the extension of the function field $\mathcal{K}$ obtained by adjoining the $\ell$-th root of $T$, and we denote this extension by $\mathcal{K}_{\ell}$. Let the polynomial ring in $\mathcal{K}_{\ell}$ be denoted by $\widetilde{\mathcal{R}}$, i.e., $\widetilde{\mathcal{R}}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\mathbb{F}_{q}[T^{1/\ell}]\subseteqq\mathcal{K}_{\ell}$. So we have the natural inclusion $\mathcal{R}\subseteqq\widetilde{\mathcal{R}}\subseteqq\mathcal{K}_{\ell}.$ It is worth noting that $\widetilde{\mathcal{R}}$ is a lattice in $\mathcal{K}_{\ell}$, while $\mathcal{R}$ is only a discrete subgroup when viewed as a subset of $\mathcal{K}_{\ell}$. We also equip $\mathcal{K}_{\ell}^{n}$ with an analogous sup norm.
@@
Given an approximating function $\psi:\{q^{r}:r\in\mathbb{Z}_{\geq 0}\}\longrightarrow\{q^{r}:r\in\mathbb{Z}\}$ and a vector $\bm{\Theta}\in\mathcal{K}^{n}$, we define the set of all $(\psi,\bm{\Theta})$-approximable vectors in $\mathcal{K}^{n}$ by
@@
$\mathcal{S}_{n}^{\bm{\Theta}}(\psi){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\left\{\mathbf{x}\in\mathcal{K}^{n}:\left\|\mathbf{x}-\frac{\mathbf{P}+\bm{\Theta}}{Q}\right\|<\frac{\psi(|Q|)}{|Q|}\quad\text{for infinitely many }(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\})\right\}.$
@@
In the special case when $\bm{\Theta}=\mathbf{0}$, the set reduces to the set of $\psi$-approximable vectors, which we denote by $\mathcal{S}_{n}({\psi}){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\mathcal{S}_{n}^{\mathbf{0}}(\psi).$
@@
Without loss of generality, we consider the manifold $\mathscr{M}$ defined by the map $\mathbf{f}:U\subseteqq\mathcal{K}^{d}\to\mathcal{K}^{n}$, where
$\mathbf{f}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}(x_{1},\ldots,x_{d},f_{1}(\mathbf{x}),\ldots,f_{m}(\mathbf{x}))=(\mathbf{x},\bm{f}(\mathbf{x})),$
@@
with $d=\dim(\mathscr{M})$, $m=\operatorname{\text{codim}}(\mathscr{M})$, $n=d+m$, and $U\subseteqq\mathcal{K}^{d}$ is an open set. We also assume that the defining map $\mathbf{f}$ is analytic (see section˜2 for the definition).
@@ 1.2. Main Results
Throughout the paper, we make the following assumptions for the sake of simplicity:
  1. ($\mathfrak{A}$.1)
    $\overline{U}$ is compact.
  2. ($\mathfrak{A}$.2)
    The map $\bm{f}=(f_{1},\dots,f_{m}):U\subseteqq\mathcal{K}^{d}\rightarrow\mathcal{K}^{m}$ admits an analytic extension to an open neighborhood of $\overline{U}$.
  3. ($\mathfrak{A}$.3)
  4. ($\mathfrak{A}$.4)
    $M>0$ is such that
@@
      $\|\bm{f}(\mathbf{x})\|\leq M$ and $\|\nabla\bm{f}(\mathbf{x})\|\leq M$, for any $\mathbf{x}\in U$; and
@@
      $|{\bar{\Phi}_{\beta}}\bm{f}(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{y}_{3})|\leq M$ for any second difference quotient $\Phi_{\beta}$ (see section˜2 for the definition) and $\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{y}_{3}\in U$.
@@
Let $n\geq 2$, let $\bm{\Theta}\in\mathcal{K}^{n}$, and let $\mathscr{M}\subseteqq\mathcal{K}^{n}$ be an analytic submanifold satisfying ($\mathfrak{A}$.1)-($\mathfrak{A}$.4). Suppose that the approximating function $\psi:\{q^{r}:r\in\mathbb{Z}_{\geq 0}\}\longrightarrow\{q^{r}:r\in\mathbb{Z}\}$ satisfies
$\sum_{\tiny{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}}\psi(|Q|)^{n}<\infty.$ (1.2)
Then almost every point of $\mathscr{M}$ is not $(\psi,\bm{\Theta})$-approximable, i.e., $\mathcal{L}_{d}\left(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi))\right)=0$.
Remark 1.3.
  1. ($\mathfrak{R}$.1)
    Given any $\varphi:\{q^{r}:r\in\mathbb{Z}_{\geq 0}\}\longrightarrow\{q^{r}:r\in\mathbb{Z}\}$, it is easy to see that
    $\sum_{\tiny{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}}\varphi(|Q|)<\infty\Longleftrightarrow\sum_{t=0}^{\infty}q^{t}\varphi(q^{t})<\infty.$ (1.3)
    In particular, (1.2) is equivalent to $\sum_{t=0}^{\infty}q^{t}\psi(q^{t})^{n}<\infty$. Indeed, given any $t\geq 0$, an easy counting argument yields that there are exactly $(q-1)q^{t}$ polynomials of degree $t$. Therefore,
    $\sum_{Q\in\mathbb{F}_{q}[T],\penalty 10000\ |Q|\leq q^{N}}\varphi(|Q|)=(q-1)\sum_{t=0}^{N}q^{t}\varphi(q^{t}),\text{ for any }N\in\mathbb{N}.$
    Note that the real analogue of this equivalence requires that $\psi$ is non-increasing (see [BY23, Lemma 1.4]), whereas here this is not necessary due to (1.3).
  2. ($\mathfrak{R}$.2)
    Fix $\kappa\in\left(1,\frac{3}{2}\right)$. It suffices to prove Theorem 1.2 for approximating functions satisfying the following:
    $\psi\left(|Q|\right)\geq\frac{1}{|Q|^{\frac{\kappa}{n}}},\penalty 10000\ \forall Q\in\mathbb{F}_{q}[T]\setminus\{0\}.$ (1.4)
    For, if $\psi$ does not satisfy (1.4), one considers the function
    $\tilde{\psi}(|Q|){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\max\left\{\psi(|Q|),\frac{1}{|Q|^{\frac{\kappa}{n}}}\right\},\penalty 10000\ \penalty 10000\ \forall Q\in\mathbb{F}_{q}[T]\setminus\{0\}.$
    It is then easy to verify that
    $\sum_{Q}\tilde{\psi}(|Q|)^{n}\leq\sum_{Q}\psi(|Q|)^{n}+\sum_{Q}\frac{1}{|Q|^{\kappa}}\asymp\sum_{t=0}^{\infty}q^{t}\psi(q^{t})^{n}+\sum_{t=0}^{\infty}q^{t}\cdot\frac{1}{q^{\kappa t}}<\infty.$
    Since $\tilde{\psi}$ clearly satisfies (1.2) and $(\psi,\bm{\Theta})$-approximability implies $(\tilde{\psi},\bm{\Theta})$-approximability, one concludes $\mathcal{L}_{d}\left(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi))\right)=0$ at once from $\mathcal{L}_{d}\left(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\tilde{\psi}))\right)=0$. In view of this, the use of the assumption mentioned above in (1.4) in the proof of Theorem 1.2 is completely justified.
A simple argument further extends Theorem 1.2 to Hausdorff measures. Recall that, for $\sigma>0$, the Hausdorff $\sigma$-measure of a set $A\subseteqq\mathcal{K}^{n}$ is defined as
$\mathcal{H}^{\sigma}(A){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\lim_{\rho\to 0^{+}}\inf\left\{\sum_{i=1}^{\infty}\operatorname{diam}(A_{i})^{\sigma}:A\subseteqq\bigcup_{i=1}^{\infty}A_{i}\text{ and }\operatorname{diam}(A_{i})<\rho,\penalty 10000\ \forall i\right\},$
where $\operatorname{diam}(A)=\sup\{\|\mathbf{u}-\mathbf{v}\|:\mathbf{u},\mathbf{v}\in A\}$ denotes the diameter of a set. The following provides an extension of Theorem 1.2 to Hausdorff measures.